# D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts

### v9

@@ IV-B The baseline for Bongard-Logo problem
We copy a baseline from Triple-CFN [32], which refers to the baseline as $f_{\theta}(z|\,x)$ that maps samples $x_{ij}$ to latent variables $z_{ij}$. The baseline employs the InfoNCE loss function, which results in the $f_{\theta}(z|\,x)$ being able to solve Bongard-logo to a certain degree. The InfoNCE loss function[41] can be mathematically expressed as:
${\ell_{\text{InfoNCE}}}({z_{pri}},{\tilde{z}_{pri}},\{{z_{au{x_{m}}}}\}_{m=1}^{M})$
$=-\log\frac{{{e^{({z_{pri}}\cdot{\tilde{z}_{pri}})/t}}}}{{{e^{({z_{pri}}\cdot{\tilde{z}_{pri}})/t}}+\sum\nolimits_{m=1}^{M}{{e^{({z_{pri}}\cdot{z_{au{x_{m}}}})/t}}}}}$ (5)
Where $z_{pri}$ and $\tilde{z}_{pri}$ represent distinct samples belonging to the representations in primary group, and $z_{au{x_{m}}}$ encompasses all $m$ samples within the auxiliary group. $m$ denotes the number of negative samples within a minimum learning unit, while $t$ represents the temperature coefficient, and $t$ is set $10^{-3}$.
This approach fosters a higher degree of similarity between $z_{pri}$ and $\tilde{z}_{pri}$ compared to the similarity between $z_{pri}$ and $z_{au{x_{m}}}$, thereby emphasizing the distinction between primary and auxiliary representations. This process aligns with the underlying logic of Bongard-logo. InfoNCE can be construed as a loss function devised for the clustering of vector groups.
This paper reproduces the baseline experiments in Triple-CFN [32], which were originally conducted on the Bongard-Logo using the ResNet18 model. The baseline focuses on learning four concepts in Bongard-Logo — ’FF’, ’BA’, ’NV’, and ’CM’ — separately or in combination [32]. To assess the impact of model size on the baseline experimental findings, we have replaced the backbone of the baseline experiments from ResNet18 to ResNet50 and repeated the experiments. The experimental results are presented in Table I, while the feedforward process is illustrated in Figure 9.
The findings indicate that when presented with a combination of databases, both ResNet18 and ResNet50 experience confusion among the four distinct human concepts. The varying degrees of decrease in the models’ reasoning accuracy on the FF and BA problems, coupled with the increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion. Furthermore, this confusion cannot be attributed to insufficient network capacity, as even ResNet50, which is more capacious than the ResNet18 used in Triple-CFN’s baseline, is subject to the same challenge.
Figure 9: Feedforward process baseline
@@ V Distribution Distance-based Discriminator of Case (D3C)
Using distributions to describe human concepts in abstract reasoning problems for the deep networks is preferable to using vectors. Distributions offer a more comprehensive and nuanced approach to capturing the diversity and uncertainty inherent in human concepts. Human concepts are often ambiguous, polysemous, and subject to varying degrees and fluctuations, which can be challenging to represent using static vectors. In contrast, distributions provide a dynamic and probabilistic framework that can accommodate this inherent variability.
By leveraging distributions to represent human concepts, we can furnish deep networks with richer and more flexible information. This enhanced representational capacity enables the networks to better understand and address the complexity and uncertainty encountered in abstract reasoning problems. For instance, probability distributions can be utilized to depict the range of possibilities and variations associated with a concept, while Gaussian distributions can model the continuity, smoothness and even discreteness of a concept’s underlying structure.
@@ V-A D3C for Bongard-Logo
D3C proposes that distributions are more suitable as abstract representations of concepts compared to vectors. Therefore, D3C decides to estimate the concept in each image of a Bongard-Logo case using a deep network in the form of $f(z|\,x)$ and represent the concept with a distribution. Finally, the distances between these distributions are measured and used as the distances between concepts. These distances can be naturally applied to solve the Bongard-Logo problem. D3C is a plug-and-play method targeted at solving Bongard-Logo and similar problems for any network in the form of $f(z|\,x)$. Notably, this brings improvements for models such as Triple-CFN and ResNet50.
In practice, this paper hypothesizes that the concepts of the image samples in a Bongard-Logo case follows a Gaussian distribution. We utilize an arbitrary backbone network $f(z|\,x)$ to excract the concepts of images within Bongard-Logo cases in form of approximating distribution. To achieve this, we modify the network to output two vectors $\mu$ and $\sigma$ instead of a single vector $z$, effectively converting $f(z|\,x)$ into $f(\mu,\sigma|\,x)$. The encoded vector $z$ is then divided into two equal-length segments $\mu$ and $\log(\sigma^{2})$, which we interpret as the mean and the logarithm of the variance, respectively. This approach normalizes the output vectors into Gaussian distributions and represents concepts as distributions in Bongard-Logo cases. Our method bears resemblance to the Variational Autoencoder (VAE) framework[8]. We argue that using distributions, rather than vectors, to describe image representations offers greater flexibility and accuracy in capturing the underlying complexity and uncertainty of abstract reasoning problems.
With the estimated concepts for each image in the Bongard-Logo cases as distributions, we compute the Sinkhorn distance between these distributions using reparameterization technique[8]. Following the logical structure of each Bongard-Logo case, we impose constraints on these distances for the backbone network $f(\mu,\sigma|\,x)$, aiming to train an effective discriminator. Specifically, we optimize the network by utilizing a binary cross-entropy loss function to enforce constraints on the exponential function of the negative Sinkhorn distance with base $e$ ($e^{-D}\in[0,1]$, where $D\in[0,+\infty]$ is the values of Sinkhorn distance), aligning with the case’s underlying logic mentioned in DC. This process can be expressed as figure 10. The D3C Method, applied to a batch of Bongard-Logo case distributions encoded by the backbone network $f(\mu,\sigma|\,x)$, can be expressed as follows:
@@
$=\sum_{i=1}^{b}\sum_{j=1}^{7}\sum_{\tilde{j}=j+1}^{7}{\text{Sinkhorn}({z_{ijs}},\,{{z}_{i\tilde{j}s}})}$
$+\sum_{i=1}^{b}\sum_{{j}=1,}^{7}\sum_{\tilde{j}=7}^{14}\log({1-{e^{-\text{Sinkhorn}({z_{ijs}},\,{{z}_{i\tilde{j}s}})}}})$ (6)
Where $b$ is the batchsize of training. The $\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\}$ represents $S$ samples drawn from the distribution $\{\mu_{ij},\sigma_{ij}|\,i\in[1,b],j\in[1,14]\}$ using the reparameterization trick, where the distribution $\{\mu_{ij},\sigma_{ij}|\,i\in[1,b],j\in[1,14]\}$ is obtained by mapping the Bongard-Logo image $\{x_{ij}|\,i\in[1,b],j\in[1,14]\}$ through the network $f(\mu,\sigma|\,x)$. The network $f(\mu,\sigma|x)$ is derived from the network $f(z|x)$ by D3C Method. $S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
@@
The design of D3C allows for its application to any image encoder that can be abstracted as $f(z|x)$. This requirement is met by a wide range of encoders, including those from the ResNet family, ViT, and specialized ones like Triple-CFN, which is particularly suitable for Bongard-Logo problems.
@@ V-B D3C on Lico-Net for RPM
Notably, in contrast to the Bongard-Logo database, both RAVEN and PGM possess well-defined and comprehensive Meta data in the form of (auxiliary) rule labels. These rule descriptions facilitate the construction of a rational concept space, with boundary concepts within the rule labels. This paper introduces the rule labels from the RPM database into the Lico-Net, aiming to enhance the reasoning accuracy and human-perspective rule interpretability of Lico-Net. This process will be implemented by the D3C method.
@@
Within the frameworks of RAVEN and PGM, there is a many-to-one mapping between image progression patterns and their corresponding rules. For instance, both single-step and double-step increments, as well as single-step and double-step decays in image attributes, are all categorized under the ‘progressive’ rule. Therefore, this paper proposes the concept of rules from the RPM database and represent them as distributions while also vectorizing the image progression patterns. This approach is precisely the original idea behind D3C, which posits that describing concepts using distributions is more appropriate than using vectors. We can leverage this approach to naturally constrain the image progression patterns within their corresponding rule distributions, thereby optimizing the deep model’s learning of rules and clarifying the boundaries among various concepts within the data.
@@
In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual descriptions of Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data. Specifically, the initial portion of the feature vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the rule descriptions and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$. Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ termed as progression pattern vectors. Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns should exhibit a high likelihood of adhering to the distribution outlined by their corresponding textual rules, and conversely, demonstrate a low probability of deviation.
A specific RPM case is encoded into $N\times M$ progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ from $L$ viewpoints by Lico-Net as aforementioned. To align Lico-Net with the D3C framework, the number of encoding patterns $N$ in Lico-Net is set according to the instance of the RPM problem, depending on the number of decoupled concepts present in the Meta data of that instance. For example, in PGM, there are two such concepts: shape pattern and line pattern, and $N$ is set to 2. The special setting of $N$ in Lico-Net can be viewed as $N$ slots, which can be aligned with the $N$ decoupled concepts described in the Meta data, respectively. By synthesizing all perspectives—specifically, by calculating the mean of the progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ across all $L$ viewpoints—Lico-Net encodes $N\times M$ vectors, denoted as $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$, representing the full-viewpoint progression patterns. This process can be expressed as:
$\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$
$=\frac{1}{L}\sum_{l=1}^{L}\{P_{nm}|\,n\in[1,N],m\in[1,M]\}_{l}$ (7)
The $l$ represents the index of viewpoints. The $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$ can be treated as $N$ sets of $M$ sampling results. This enables the optimization of distribution distances based on the Sinkhorn distance, thereby enhancing the overall performance through D3C framework. Apparently, this paper utilizes the Sinkhorn distance to constrain the distance between $M$ progression pattern vectors and their corresponding distributions’ sampling results. The $N$ decoupling concept presented in Meta data is employed to concurrently impose constraints on the $N$ sets of $M$ progression pattern vectors, and the process of aligning $\{\overline{P}_{nm}\}$ with the distributional concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ can be illustrated in figure 11.
@@
$=\sum_{n=1,\tilde{k}|\,n}^{N}{\text{Sinkhorn}({\overline{P}_{nm}},\,{r_{\tilde{k}s}})}$
$+\sum_{k=1,k\neq\tilde{k}}^{K}\sum_{n=1,\tilde{k}|\,n}^{N}\log({1-{e^{-\text{Sinkhorn}({\overline{P}_{nm}},\,{r_{{k}s}})}}})$ (8)
The $\overline{P}_{nm}$ represents the full-viewpoint progressive pattern vectors encoded by Lico-Net, and $\{r_{ks}\}$ represents the $S$ sampling results from the distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of the rule description. It is evident that the specific value of $\tilde{k}$ is contingent upon the description of Meta data corresponding to the particular case of the RPM problem. symbol “$\tilde{k}|\,n$” represents that the specific value of $\tilde{k}$ is depended on $n$. This dependency realizes the fact that each slots in Lico-Net are aligned with their corresponding concepts. The $\text{Sinkhorn}(\cdot,\cdot)$ represents the Sinkhorn distance operator. Other logic of symbol in function (V-B) has mentioned. It becomes evident that we are employing the Binary Cross Entropy (BCE) loss function to impose constraints on the exponential negative Sinkhorn distance result. The feedforward process is illustrated in the figure 12.
Figure 12: Feedforward process of Lico-Net aligned with D3C
@@ V-C A low-cost loss function for the D3C, the Hermes for the Zeus
Sampling $S$ times (where $S$ is arbitrarily set, and $S$ is set to $20$ for Sinkhorn distance) from the constructed distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of concepts in Meta data and yielding $\{r_{ks}|\,k\in[1,K],s\in[1,S]\}$, we calculate the Sinkhorn distance between the progression pattern vectors and these sampling results $\{r_{ks}\}$. This Sinkhorn distance approximates the Wasserstein distance, which is currently a linear programming algorithm also known as the Earth Mover’s Distance (EMD) algorithm. Typically, this approach requires a substantial number of samples and incurs a considerable computational cost. Even if we treat the $M$ progression pattern vectors as $M$ sampling results, it is still difficult to satisfy the prerequisite of extensive sampling in the context of the RPM problem, where $M$ is set to 9 in PGM and 3 in RAVEN. To reduce computational cost and sampling frequency, this paper employs cosine similarity between vectors as an alternative distance metric norm to the Sinkhorn distance. Obviously, the cosine similarity satisfacte with the $Lipschitz$ continuity condition. For the inner product, treating it as a function with one fixed vector, we observe that it exhibits $Lipschitz$ continuity. Specifically, given a fixed vector $v$, the function $f(u)=u\cdot v$ is $Lipschitz$ continuous with respect to $u$, with the $Lipschitz$ constant $L$ being equivalent to the norm of $v$. This property is attributed to the inner product’s adherence to the Cauchy-Schwarz inequality: $|u\cdot v|\leq||u||\cdot||v||$. Leveraging this inequality, we are able to deduce the underlying inequality:
$f(u_{1})-f(u_{2})=|u_{1}\cdot v-u_{2}\cdot v|=|(u_{1}-u_{2})\cdot v|\leq||u_{1}-u_{2}||\cdot||v||$ (9)
In our study, we opted for cosine similarity, considering normalized vectors $u$ and $v$, where $||v||$ and $||u||$ are equal to 1, and Lipschitz constant $L=||v||=1$. It is reasonable to choose the mean of cosine similarity between sampling vectors as the distance metric function instead of Sinkhorn distance. The calculate process of new method applied on Bongard-Logo can be expressed as:
@@
From a certain perspective, estimating the distribution distance based on cosine similarity between sampling results can be regarded as a variant of the Wasserstein distance based on cosine similarity that relinquishes the lower bound estimation and possesses a relatively low, fixed Lipschitz constant. Therefore, this method, we called “D3C-cos”, cannot precisely quantify the distance between distributions. It serves as a low-cost loss function for constrained inter-distribution distance, serving as a surrogate for the Sinkhorn distance, the Hermes for the Zeus.
The success of such simplification can be attributed, in part, to our modeling of the distribution of pattern descriptions as a Gaussian distribution.

### v12

@@ IV-B The baseline for Bongard-Logo problem
We copy a baseline from Triple-CFN [32], which refers to the baseline as $f_{\theta}(z|\,x)$ that maps samples $x_{ij}$ to latent variables $z_{ij}$. The baseline employs the InfoNCE loss function, which results in the $f_{\theta}(z|\,x)$ being able to solve Bongard-logo to a certain degree. This approach fosters a higher degree of similarity between $z_{ij}$ and $z_{i\tilde{j}}$ (where $\tilde{j}\neq j$ and $\tilde{j}$, $j$ $\in[1,7]$), compared to the similarity between $z_{ij}$ and $\{z_{ij}|\,j\in[8,13]\}$, thereby emphasizing the distinction between primary and auxiliary representations. This process aligns with the underlying logic of Bongard-logo.
This paper reproduces the baseline experiments in Triple-CFN [32], which were originally conducted on the Bongard-Logo using the ResNet18 model. The baseline focuses on learning four concepts in Bongard-Logo — ‘FF’, ‘BA’, ‘NV’, and ‘CM’ — separately or in combination [32]. To assess the impact of model size on the baseline experimental findings, we have replaced the backbone of the baseline experiments from ResNet18 to ResNet50 and repeated the experiments. The experimental results are presented in Table I.
The findings indicate that when presented with a combination of databases, both ResNet18 and ResNet50 experience confusion among the four distinct human concepts. The varying degrees of decrease in the models’ reasoning accuracy on the FF and BA problems, accompanied by an increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion. Furthermore, this confusion cannot be attributed to insufficient network capacity, as even ResNet50, which is more capacious than the ResNet18 used in Triple-CFN’s baseline, is subject to the same challenge.
@@ V Distribution Distance-based Discriminator of Case (D3C)
Using distributions to describe human concepts in abstract reasoning problems for deep neural networks is advantageous compared to the use of vectors. This is because distributions present a more comprehensive and nuanced means of capturing the inherent diversity and uncertainty in human concepts. Human concepts are frequently characterized by ambiguity, polysemy, intersectionality, and overlap, making their representation via static vectors challenging. Conversely, distributions afford a dynamic and probabilistic framework that can effectively accommodate such inherent variability. Employing distributions to represent human concepts endows deep neural networks with richer and more flexible information. This, in turn, enhances their capability to comprehend and resolve complex, uncertain abstract reasoning problems. Probability distributions illustrate the possibilities and variations associated with a concept, whereas Gaussian distributions specifically model the underlying structure of a concept.
@@ V-A D3C for Bongard-Logo
D3C argues that distributions are better abstract concepts than vectors. Therefore, it employs a deep network $f(z|\,x)$ to estimate concepts in Bongard-Logo images and represents them as distributions. By measuring concept distances as the distances between these distributions, the method becomes applicable to solving the Bongard-Logo problem. D3C is a plug-and-play method for any network with form of $f(z|\,x)$, improving models like Triple-CFN and ResNet50.
In practice, this paper assumes the concepts in Bongard-Logo cases follow a Gaussian distribution. We use a backbone network $f(z|\,x)$ to extract these concepts as distributions. To do this, we adapt the network to output $\mu$ and $\sigma$ instead of a single vector $z$, turning $f(z|\,x)$ into $f(\mu,\sigma|\,x)$. The vector $z$ is split into two equal-length segments $\mu$ and $\log(\sigma^{2})$, representing mean and variance logarithm, respectively. This method normalizes original outputs $z$ to Gaussian distributions $\{\mu,\sigma^{2}\}$ and bears resemblance to the Variational Autoencoder (VAE) framework[8]. Representing concepts as distributions, not vectors, provides more flexibility and accuracy in abstract reasoning.
With the estimated concepts for each image in the Bongard-Logo dataset represented as distributions, we compute the Sinkhorn distance between these distributions using the reparameterization technique[8]. Subsequently, we impose constraints on these distances to train and optimize the backbone network $f(z|\,x)$, aligning with the underlying logic of the cases. The optimization process involves minimizing a binary cross-entropy loss, which is calculated using the exponential of the negative Sinkhorn distance as a factor. The entirety of this process is illustrated in Figure 8. When applied to a batch of Bongard-Logo cases, the D3C method can be summarized as follows:
@@
$=\sum_{i=1}^{b}\sum_{j=1}^{7}\sum_{\tilde{j}=j+1}^{7}{\text{S}(\{z_{ijs}|\,s\in[1,S]\},\,\{z_{i\tilde{j}s}|\,s\in[1,S]\})}$
$+\sum_{i=1}^{b}\sum_{{j}=1,}^{7}\sum_{\tilde{j}=7}^{14}\log({1-{e^{-\text{S}(\{z_{ijs}|\,s\in[1,S]\},\,\{z_{i\tilde{j}s}|\,s\in[1,S]\})}}})$ (2)
[…]
@@
The design of D3C allows it to be applied to any image encoder that can be abstracted as $f(z|x)$. A wide range of encoders, including those from the ResNet family, ViT, and specialized encoders like Triple-CFN—which is particularly suitable for Bongard-Logo problems—meet the abstraction requirement and are thus compatible with D3C.
@@ V-B D3C on Lico-Net for RPM
Notably, in contrast to the Bongard-Logo database, each case in RAVEN and PGM is accompanied by well-defined and comprehensive rule labels in the form of Meta data. These auxiliary labels provide descriptions of the image progression patterns within the cases, facilitating the delineation of concept boundaries. Therefore, this paper decides to introduce the Meta data from the RPM database into the Lico-Net, aiming to organize the distribution of the encoded progressive pattern vector $\{P_{nm}\}$ in Lico-Net. This enhancement is expected to improve both the reasoning accuracy and human-perspective interpretability of Lico-Net. This process will be implemented by the D3C method.
@@
Within the frameworks of RAVEN and PGM, there exists a many-to-one mapping between image progression patterns and their corresponding concepts. For instance, both single-step and double-step increments, as well as single-step and double-step decrements in image attributes, are categorized under the umbrella of the “progressive” concept according to Meta data. Therefore, this paper proposes extracting the concepts described by the Meta data and representing them as distributions, which aligns with the original idea behind D3C. D3C posits that using distributions to describe concepts is more suitable than using vectors. As Lico-Net has already extracted the progression patterns of images for the RPM matrix and encoded them as vectors, we can naturally employ concepts with distribution forms to constrain and delimit these image progression pattern vectors, ensuring that these progression pattern vectors belong to the distribution of corresponding concept. This approach will optimize Lico-Net’s learning of rules and clarify the boundaries among various concepts. The aforementioned idea can be represented in the framework illustrated in the Figure 9.
Figure 9: The framework of D3C
@@
In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data. Specifically, the initial portion of the concept vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the concepts described in Meta data and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$. Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ termed as progression pattern vectors. Lico-Net extracts N types of progression patterns (where N can be freely set) for each complete RPM matrix, and encodes M progression pattern vectors (where M is determined by the number of existing minimal concept groups) for each progression pattern, forming the vector group $\{P_{nm}\}$. This process can be reviewed in Figure 6. To integrate Lico-Net with the D3C framework, the number of encoding patterns $N$ in Lico-Net, which can no longer be set arbitrarily, is set according to the instance of the RPM problem, depending on the number of decoupled concepts present in the Meta data of that instance. For example, in PGM, there are two such concepts: shape pattern and line pattern, and $N$ is set to 2. The special setting of $N$ in Lico-Net can be viewed as $N$ slots, which can be aligned with the $N$ decoupled concepts described in the Meta data, respectively. Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns $\{P_{nm}\}$ should exhibit a high likelihood of adhering to the distribution outlined by their corresponding concepts, and conversely, demonstrate a low probability of deviation.
To impose constraints on the distributions of the extracted $N$ progression patterns, we have devised the following process. By synthesizing all perspectives—specifically, by calculating the mean of the progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ across all $L$ viewpoints—Lico-Net encodes $N\times M$ vectors, denoted as $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$, representing the full-viewpoint progression patterns. The set $\{\overline{P}_{nm}\}$ can be interpreted as the outcome of sampling $M$ times from each of the $N$ distributions of progression patterns across all viewpoints generated by Lico-Net. This interpreting perspective enables us to utilize Meta data to enforce constraints on the $N$ distributions. Next, we sample $S$ times from the distribution of concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ to obtain sampling results $\{r_{ks}|\,k\in[1,K],s\in[1,S]\}$. Consequently, this paper employs the Sinkhorn distance to constrain the distance between the distributions of the $N$ kinds of progression patterns encoded by Lico-Net and their corresponding concept distributions in the Meta data. The process of aligning $\{\overline{P}_{nm}\}$ with the distribution of concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ can be illustrated in figure 10.
@@
$=\sum_{n=1,\tilde{k}|\,n}^{N}{\text{S}(\{\overline{P}_{nm}|\,m\in[1,M]\},\,\{r_{\tilde{k}s}|\,s\in[1,S]\})}$
$+\sum_{k=1,k\neq\tilde{k},\tilde{k}|\,n}^{K}\sum_{n=1}^{N}\log({1-{e^{-\text{S}((\{\overline{P}_{nm}|\,m\in[1,M]\},\,\{r_{{k}s}|\,s\in[1,S]\}))}}})$ (3)
[…]
In summary, Lico-Net is capable of extracting $N$ types of progression patterns for each minimal concept group in the RPM problem. Given the existence of $M$ minimal concept groups, we obtain a total of $N\times M$ progression pattern vectors. To constrain and clarify the distribution of vectors under each of the $N$ progression patterns, we consider the vectors as the outcome of $M$ samplings from each of the $N$ distributions. This perspective enables us to implement distributional constraints based on concepts in Meta data, thereby enhancing the alignment and interpretability of the progression patterns extracted by Lico-Net.
@@ V-C A low-cost loss function for the D3C, the Hermes for the Zeus
Sampling $S$ times (where $S$ is arbitrarily set, and $S$ is set to $20$ for Sinkhorn distance) from the constructed distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of concepts in Meta data and yielding $\{r_{ks}|\,k\in[1,K],s\in[1,S]\}$, we calculate the Sinkhorn distance between the progression pattern vectors and these sampling results $\{r_{ks}\}$. Typically, this approach requires a substantial number of samples and incurs a considerable computational cost. To reduce computational cost and sampling frequency, this paper employs the mean of cosine similarity between sampling vectors as an alternative distance metric norm to the Sinkhorn distance.
The calculate process of new method applied on Bongard-Logo can be expressed as:
@@
Estimating distribution distance via cosine similarity between samples can be seen as a modified Wasserstein distance, relinquishing lower bound estimation and having a fixed Lipschitz constant. This ”D3C-cos” method, while not exact, provides a cost-effective loss function for distribution distance, acting as a substitute for the Sinkhorn distance. We still use Sinkhorn distance for performance testing. The success of such simplification can be attributed, in part, to our modeling of the distribution of pattern descriptions as a Gaussian distribution.